OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify the square root of 180.and provide the simplified form. You can show a square root by writing "sqrt()". For example, the square root of 2 can be written as "sqrt(2)". Can someone explain a method i could slove this?

OpenStudy (sonyalee77):

The best way to determine the square root of an integer is to first determine its prime factors. Thus 180 = 2 x 2 x 3 x 3 x 5 = 2^2 x 3^2 x 5 Hence sqrt(180) = sqrt(2^2 x 3^2 x 5) = 2 x 3 x sqrt(5) = 6*sqrt(5) = But wait a mo. A number x has two square roots; So 180 -> +/- sqrt(180) = +/- 6*sqrt(5) For me, that answer is complete as sqrt(5), being an irrational number cannot be accurately described in a fixed number of decimal of places. However, if you wish to express the square root(s) of 180 to 3 decimal places say, then the solutions are +/- sqrt(180)= +/- 13,416

OpenStudy (sonyalee77):

is this good enough?

OpenStudy (anonymous):

Oh Well that does seem to make sense.

OpenStudy (anonymous):

Thank you for the explaination Sonya

OpenStudy (sonyalee77):

No problem :)